# Bias Analysis and Mitigation through Protected Attribute Detection and Regard Classification

### v1

@@ 1 Introduction
[…]
@@ 2.1 Protected Attribute Detection
While traditional WSD relies on glosses defined in WordNet Miller (1995), this is not suitable for our purpose, e.g. since there are no sense inventories which distinguish the usages of “asian” in Table 1. Therefore, for each of the 97 keywords, we crafted simple yet sufficient glosses describing the protected attributes. [^2] It is worth noting that our glosses are NOT intended to define strict word senses but to practically distinguish whether the keyword indicates a protected *human* attribute. We show several illustrative examples in Table 3 (Appendix A.1).
@@ 2.2 Regard Classification
Regard classification is widely used to evaluate the bias of LLM generated texts (Sheng et al., 2019; Dhamala et al., 2021). In this study, we apply this technique to pretraining data analysis.
Unfortunately, existing regard classifier (Sheng et al., 2019) is only trained on templated texts (e.g. “XYZ was regarded as …”) which makes it difficult to be applied on realistic texts with diverse syntactic structures. Therefore, similar to our WSD (Section 2.1), we generated a new training data for regard classification based on an existing LLM.
@@ 2.3 Annotation Agreement
Regarding WSD for protected attribute detection (Section 2.1), we computed the annotation agreement between Mixtral-8x22B and Gloss-RoBERTa based on Cohen’s kappa. On average across all attributes, we observed a moderate agreement of around 0.59. Additionally, for a small set of attributes, we conducted manual annotation and confirmed that agreement (1) among human annotators and (2) between humans and these models are also at similar levels (0.56 to 0.70). Therefore, our models can disambiguate protected attributes at near human-like levels (see Table 5 for example predictions).
As for regard classification, we prepared a high-quality test set double-checked with Mixtral-8x7B and 8x22B. Based on our RoBERTa-based classifier, we achieved a decent F1 performance of 0.91 on micro and 0.82 on macro average across all attributes. This verifies that we could successfully distill the regard judgements of Mixtral into a much more light-weight classifier (cf. Table 6).
@@ 3 Experiments
As a preparation, we first sampled over 3M English documents from CC and extracted sentences of appropriate length. [^3] Specifically, more than 16 tokens and less than 128 tokens based on the NLTK tokenizer Bird (2006). Then, we conducted protected attribute detection (Section 2.1) and regard classification (Section 2.2) on these sentences. In the following experiments, we used up to 100K sentences per attribute to demonstrate the effect of our bias analysis (Section 3.1) and mitigation (Section 3.2) techniques.
@@ 3.1 Bias Analysis
Here, the left term represents frequency bias (eq. (1)) and the right term regard bias of when a word $w$ and attribute $a$ co-occur. For instance, if $w$ = supremacist and $a$ = white tend to co-occur with negative regard, the score increases for $r$ = negative. This way, we can analyze not only whether $w$ and $a$ co-occur, but whether they co-occur accompanying specific regard in the pretraining corpora.
@@ 3.2 Bias Mitigation
In this study, we verify that problematic word association can be alleviated by properly balancing the regard distributions. Specifically, we ensure the negative regard ratio to be at most 1% for all protected attributes by downsampling negative regard sentences. Then, we compute the relative reduction in the conditional probability $p(w|a)$, where $w$ is a negatively biased word towards $a$ [^4] To be precise, we report $100\!\times\!\frac{p^{\prime}(w|a)}{p(w|a)}$%, where $p$ and $p^{\prime}$ are probabilities before and after intervention, respectively.
Through this intervention, we confirmed $p(w|a)$ can be dramatically reduced, e.g. down to 19% and 18% for $w=$supremacist” and “racists” toward $a=$white”, respectively. Similarly, $p(w|a)$ can be diminished to 46% and 26% for $w=$terrorist” and “assaults” toward $a=$arab”, while neutral or positive word associations remain unchanged or slightly increase, e.g. up to 112% for $w=$generosity” and “hospitality”.
@@ 4 Discussion and Conclusion
While our taxonomy of protected attributes cover a reasonably wide range of demographics, it does not currently support attribute classes such as age, marital status, and political belief, nor our list of attributes in each class is comprehensive, e.g. many attributes are still missing in the class of nationality, disability, etc. Expanding and refining the taxonomy is a subject of future study, which we expect can be approached by leveraging LLMs and human-curated resources (Smith et al., 2022).
@@
In terms of regard classification, our current classifier is distilled from (i.e. relies on the judgements of) Mixtral-8x7B-Instruct-v0.1, which themselves may be biased in some undesirable ways. While we expect this issue can be alleviated by collecting multiple judgements from a diverse pool of human annotators or LLMs, developing a more reliable regard classier remains as an important future work.
As for bias mitigation, we’ve obtained promising results on reducing negative word associations through regard balancing: however, we’ve not conducted a full ablation study of pretraining LLMs from scratch on the regard balanced datasets. Due to the extreme cost of the above experiments, we leave it as an opportunity for future study.
@@ A.2 Mixtral Prompts
WSD for Protected Attribute Detection (Section 2.1) Below is a text including the keyword “{Keyword}”. Determine whether the keyword indicates a person (or people) {Gloss}. - Focus on the keyword “{Keyword}”. Briefly explain its usage in the text and then answer with yes, no, or unsure. - Answer yes if there is a clear reference to the person (or people) {Gloss}. - Answer no if there is no such reference, e.g. the keyword describes *non-human* entities (e.g. animals, objects, events) or does not indicate the individual(s) {Gloss}. - Answer unsure otherwise, e.g. if the reference is unclear, indirect, or implicit. Text: {Text} After examining the text: - Briefly justify your answer up to 100 words, and then answer with yes, no, or unsure. - Conclude your response in the format: “Therefore, the answer is …”. Regard Classification (Section 2.2) What best describes the regard towards the person (or people) {Gloss} in the following text? Focus on the individual(s) indicated by the keyword “{Keyword}” and answer from the following options: - Positive: The regard for the individual(s) in the text is predominantly positive, e.g. containing descriptors related to positive personality, high capability, significant achievements, etc. - Negative: The regard for the individual(s) in the text is predominantly negative, e.g. containing descriptors related to negative personality, lack of ability, commitment of illegal activities, etc. - Neutral: There is no clear impact of regard for the individual(s) associated with the text. Text: {Text} After examining the text: - Start your response with the answer (positive, negative, or neutral), followed by justification up to 100 words.
Table 4: Mixtral prompts used in word sense disambiguation (WSD) for protected attribute detection (Section 2.1) and regard classification (Section 2.2). Similar to the input format in Table 1, {Text}  indicates the text containing the keyword, {Keyword}  the target keyword, and {Gloss}  the keyword’s defined gloss.
@@
Table 5: Illustrative examples of word sense disambiguation (WSD) for protected attribute detection (Section 2.1) where both Mixtral-8x22B and Gloss-RoBERTa agreed on the predictions.
@@
Table 6: Illustrative examples of regard classification towards protected attributes (Section 2.2) where both Mixtral-8x7B and our RoBERTa-based classifier agreed on the predictions.
@@
Regarding WSD for protected attribute detection (Section 2.1), we applied chain-of-thought (CoT) prompting (Wei et al., 2022) to first explain the keyword usage in the context, which consistently improved the prediction quality. Based on Mixtral’s response, we considered the answer of yes as protected attribute and unsure/no as non-protected attribute. In Table 5, we show several examples of WSD results where both Mixtral-8x22B and Gloss-RoBERTa agreed on the predictions.
As for regard classification (Section 2.2), we generally followed the criteria of Sheng et al. (2019) and explicitly included them in the prompt. In Table 6, we show several illustrative examples where both Mixtral-8x7B and our RoBERTa-based classifier agreed on the predictions.
@@
Recall@ Frequency Bias Frequency+Regard Bias Positive Stereotypes (mean offensiveness score $=-1$) 50 2.67 5.01 100 4.00 7.79 200 7.23 13.24 300 9.79 16.80 500 15.80 24.69 Negative Stereotypes (mean offensiveness score $\geq 1$) 50 3.68 13.97 100 9.56 17.65 200 13.97 33.09 300 16.18 38.97 500 25.00 47.06
@@ Appendix B Bias Analysis Details
In Table 7, we show (partial) results of our bias analysis (Section 3.1) for the class $A=$ nationality. Due to the large number of attributes in this class, we restricted our analysis to 53 nationalities which are covered in the SeeGULL dataset (Jha et al., 2023). [^6] https://github.com/google-research-datasets/seegull. SeeGULL is a high quality and broad coverage resource containing prevalent stereotypes, which we use for the verification of our analysis.
Specifically, we utilize their offensiveness score annotations and consider the stereotype in their list as positive if the mean score is --1 (e.g. “wise” for chinese) and negative if it is greater than or equal to 1 (e.g. “greedy” for japanese). Considering these stereotypes as the ground truths, we measured the recall@k for the 53 nationalities based on the vocabulary sorted by the frequency bias and frequency+regard bias scores. [^7] As for the frequency+regard bias scores, we set the regard $r$ = positive/negative in eq. (2) to compute the alignment with SeeGULL’s positive/negative stereotypes, respectively.
In Table 8, we show the results of the alignment with SeeGULL stereotypes. From these results, we can verify that frequency+regard bias scores can identify positive/negative stereotypes more accurately, as (partially) illustrated in Table 7. In fact, our analysis can detect potential stereotypes that may be even missing in SeeGULL (e.g. “technological” or “racist” against chinese people), demonstrating the effectiveness of our approach for comprehensively auditing social biases.

### v2

@@ 1 Introduction
Input Format: [BOS] {Text}[SEP] {Keyword}; {Gloss}[EOS] Prediction [BOS] I observed a group of Asian visitors … [SEP] asian; a person of asian race/ethnicity [EOS] ✓ Protected Attribute [BOS] … and more traditional Asian cuisine. [SEP] asian; a person of asian race/ethnicity [EOS] $\times$  Non-Protected Attribute
@@ 2.1 Protected Attribute Detection
While traditional WSD relies on glosses defined in WordNet Miller (1995), this is not suitable for our purpose, e.g. since there are no sense inventories which distinguish the usages of “asian” in Table 1. Therefore, for each of the 97 keywords, we crafted simple yet sufficient glosses describing the protected attributes. [^2] It is worth noting that our glosses are NOT intended to define strict word senses but to practically distinguish whether the keyword indicates a protected human attribute. We show several illustrative examples in Table 3 (Appendix A.1).
@@ 2.2 Regard Classification
Regard classification is widely used to evaluate the bias of LLM generated texts (Sheng et al., 2019; Dhamala et al., 2021). In this study, we aim to apply this technique to pretraining data analysis.
Unfortunately, existing regard classifier (Sheng et al., 2019) is only trained on templated texts (e.g. “XYZ was regarded as …”) which makes it difficult to be applied on realistic texts with diverse syntactic structures. Therefore, similar to our WSD (Section˜2.1), we generated new training data for regard classification based on an existing LLM.
@@ 2.3 Annotation Agreement
Regarding WSD for protected attribute detection (Section˜2.1), we computed the annotation agreement between Mixtral-8x22B and Gloss-RoBERTa based on Cohen’s kappa. On average across all attributes, we observed a moderate agreement of around 0.59. Additionally, for a small set of attributes, we conducted manual annotation and confirmed that agreement (1) among human annotators and (2) between humans and these models are also at similar levels (0.56 to 0.70). Therefore, our models can disambiguate protected attributes at near human-like levels (see Table 5 for example predictions).
As for regard classification, we prepared a high-quality test set double-checked with both Mixtral-8x7B and 8x22B for consistency. Based on our RoBERTa-based classifier, we achieved a decent F1 performance of 0.91 on micro and 0.82 on macro average across all attributes. This verifies that we could successfully distill the regard judgements of the strong Mixtral teacher into a much more light-weight classifier (cf. Table 6).
@@ 3 Experiments
As a preparation, we first sampled over 3M English documents from CC and extracted sentences of appropriate length. [^5] Specifically, more than 16 tokens and less than 128 tokens based on the NLTK tokenizer Bird (2006). Then, we conducted protected attribute detection (Section˜2.1) and regard classification (Section˜2.2) on these sentences. In the following experiments, we used up to 100K sentences per attribute to demonstrate the effect of our bias analysis (Section˜3.1) and mitigation (Section˜3.2) techniques.
@@ 3.1 Bias Analysis
Here, the left term represents frequency bias (eq. (1)) and the right term regard bias of when a word $w$ and attribute $a$ co-occur. [^6] While we ignore the difference in the scale of these bias scores, this can be easily adjusted (e.g. by introducing scaling factors) if necessary. For instance, if $w$ = supremacist and $a$ = white tend to co-occur with negative regard, the score increases for $r$ = negative. This way, we can analyze not only whether $w$ and $a$ co-occur, but whether they co-occur accompanying specific regard in the pretraining corpora.
@@ 3.2 Bias Mitigation
In this study, we verify that problematic word association can be alleviated by properly balancing the regard distributions. Specifically, we ensure the negative regard ratio to be at most 1% for all protected attributes by downsampling negative regard sentences. Then, we compute the relative reduction in the conditional probability $p(w|a)$, where $w$ is a negatively biased word towards aa. [^7] To be precise, we report $100\!\times\!\frac{p^{\prime}(w|a)}{p(w|a)}$%, where $p$ and $p^{\prime}$ are probabilities before and after intervention, respectively.
Through this intervention, we confirmed $p(w|a)$ can be dramatically reduced, e.g. down to 19% and 18% for w=w=supremacist” and “racists” toward a=a=white”, respectively. Similarly, $p(w|a)$ can be diminished to 46% and 26% for w=w=terrorist” and “assaults” toward a=a=arab”, while neutral or positive word associations remain unchanged or slightly increase, e.g. up to 112% for w=w=generosity” and “hospitality”.
@@ 4 Discussion and Conclusion
While our taxonomy of protected attributes covers a reasonably wide range of demographics, it does not currently support attribute classes such as age, marital status, and political belief, nor our list of attributes in each class is comprehensive, e.g. many attributes are still missing in the class of nationality, disability, etc. Expanding and refining the taxonomy is a subject of future study, which we expect can be approached by leveraging LLMs and human-curated resources (Smith et al., 2022).
@@
In terms of regard classification, our current classifier is distilled from (i.e. relies on the judgements of) Mixtral-8x7B-Instruct, which themselves may be biased in some undesirable ways. While we expect this issue can be alleviated by collecting multiple judgements from a diverse pool of human annotators or LLMs, developing a more reliable regard classier remains an important future work.
As for bias mitigation, we obtained promising results on reducing negative word associations through regard balancing: however, we’ve not conducted a full ablation study of pretraining LLMs from scratch on the regard balanced datasets. Due to the extreme cost of the above experiments, we leave it as an opportunity for future study.
@@ A.2 Mixtral Prompts
WSD for Protected Attribute Detection (Section˜2.1) Below is a text including the keyword “{Keyword}”. Determine whether the keyword indicates a person (or people) {Gloss}. - Focus on the keyword “{Keyword}”. Briefly explain its usage in the text and then answer with yes, no, or unsure. - Answer yes if there is a clear reference to the person (or people) {Gloss}. - Answer no if there is no such reference, e.g. the keyword describes *non-human* entities (e.g. animals, objects, events) or does not indicate the individual(s) {Gloss}. - Answer unsure otherwise, e.g. if the reference is unclear, indirect, or implicit. Text: {Text} After examining the text: - Briefly justify your answer up to 100 words, and then answer with yes, no, or unsure. - Conclude your response in the format: “Therefore, the answer is …”. Regard Classification (Section˜2.2) What best describes the regard towards the person (or people) {Gloss} in the following text? Focus on the individual(s) indicated by the keyword “{Keyword}” and answer from the following options: - Positive: The regard for the individual(s) in the text is predominantly positive, e.g. containing descriptors related to positive personality, high capability, significant achievements, etc. - Negative: The regard for the individual(s) in the text is predominantly negative, e.g. containing descriptors related to negative personality, lack of ability, commitment of illegal activities, etc. - Neutral: There is no clear impact of regard for the individual(s) associated with the text. Text: {Text} After examining the text: - Start your response with the answer (positive, negative, or neutral), followed by justification up to 100 words.
Table 4: Mixtral prompts used in word sense disambiguation (WSD) for protected attribute detection (Section˜2.1) and regard classification (Section˜2.2). Similar to the input format in Table 1, {Text}  indicates the text containing the keyword, {Keyword}  the target keyword, and {Gloss}  the keyword’s defined gloss.
@@
Table 5: Illustrative examples of word sense disambiguation (WSD) for protected attribute detection (Section˜2.1) where both Mixtral-8x22B and Gloss-RoBERTa agreed on the predictions.
@@
Table 6: Illustrative examples of regard classification towards protected attributes (Section˜2.2) where both Mixtral-8x7B and our RoBERTa-based classifier agreed on the predictions.
@@
Regarding WSD for protected attribute detection (Section˜2.1), we applied chain-of-thought (CoT) prompting (Wei et al., 2022) to first explain the keyword usage in the context, which consistently improved the prediction quality. Based on Mixtral’s response, we considered the answer of yes as protected attribute and unsure/no as non-protected attribute. In Table 5, we show several examples of WSD results where both Mixtral-8x22B and Gloss-RoBERTa agreed on the predictions.
As for regard classification (Section˜2.2), we generally followed the criteria of Sheng et al. (2019) and explicitly included them in the prompt. In Table 6, we show several illustrative examples where both Mixtral-8x7B and our RoBERTa-based classifier agreed on the predictions.
@@
Recall@ Frequency Bias Frequency+Regard Bias Positive Stereotypes (mean offensiveness score =1=-1) 50 2.67 5.01 100 4.00 7.79 200 7.23 13.24 300 9.79 16.80 500 15.80 24.69 Negative Stereotypes (mean offensiveness score $\geq 1$) 50 3.68 13.97 100 9.56 17.65 200 13.97 33.09 300 16.18 38.97 500 25.00 47.06
@@ Appendix B Bias Analysis Details
In Table 7, we show (partial) results of our bias analysis (Section˜3.1) for the class $A=$ nationality. Due to the large number of attributes in this class, we restricted our analysis to 53 nationalities which are covered in the SeeGULL dataset (Jha et al., 2023). [^9] https://github.com/google-research-datasets/seegull. SeeGULL is a high quality and broad coverage resource containing prevalent stereotypes, which we use for the verification of our analysis.
Specifically, we utilize their offensiveness score annotations and consider the stereotype in their list as positive if the mean score is -1 (e.g. “wise” for chinese) and negative if it is greater than or equal to 1 (e.g. “greedy” for japanese). Considering these stereotypes as the ground truths, we measured the recall@k for the 53 nationalities based on the vocabulary sorted by the frequency bias and frequency+regard bias scores. [^10] As for the frequency+regard bias scores, we set the regard $r$ = positive/negative in eq. (2) to compute the alignment with SeeGULL’s positive/negative stereotypes, respectively.
In Table 8, we show the results of the alignment with SeeGULL stereotypes. From these results, we can verify that frequency+regard bias scores can identify positive/negative stereotypes more accurately, as (partially) illustrated in Table 7. In fact, our analysis can detect potential stereotypes that may be even missing in SeeGULL (e.g. “technological” or “racist” against chinese people), demonstrating the effectiveness of our approach.